OpenStudy (anonymous):

If x^2−14x+1=0, what is the value of x+1/x?

OpenStudy (anonymous):

\(\frac{x+1}{x}\)?

OpenStudy (anonymous):

no \[x + \frac{ 1 }{ x }\]

OpenStudy (anonymous):

@satellite73

OpenStudy (shubhamsrg):

you have to find (x^2 +1)/x from the given equation, you can easily see what x^2 +1 is rest is a piece of cake.

OpenStudy (anonymous):

huh?

OpenStudy (shubhamsrg):

x^2 - 14x + 1 =0 so x^2 +1 = ?

OpenStudy (anonymous):

14x?

OpenStudy (shubhamsrg):

right and you have to find the value of (x^2 + 1)/x

OpenStudy (anonymous):

x = 14, 13 right?

OpenStudy (shubhamsrg):

14,13 ? what ?

OpenStudy (anonymous):

oh...i remember now...the sum of the roots and its reciprocal of a quadratic equation is just -b

OpenStudy (anonymous):

thus -b = 14

OpenStudy (shubhamsrg):

thats not true. Thats only true when a=c=1, which was your case here.

OpenStudy (shubhamsrg):

not 1, its true when a=c.

OpenStudy (anonymous):

1) find the zeros, \(x_1\) and \(x_2\) 2) compute \(x_1+\frac{1}{x_1}\)

OpenStudy (anonymous):

oh....what if \[a \neq c \]?

OpenStudy (shubhamsrg):

if a not equals c, then thats not true then you might need to find the roots separately , or see if anything alternative is applicable. In this case, you don;t need to find the respective roots.